OpenStudy (ifrah123):

let (a,57)=1 then a^18 congruent to 1 in modulo 57.

ganeshie8 (ganeshie8):

Little Fermat gives you a^18 = 1 ( mod 19 ) a^2 = 1 ( mod 3 )

ganeshie8 (ganeshie8):

second congruence implies a^18 = 1 ( mod 3 )

ganeshie8 (ganeshie8):

That means a^18 - 1 is divisible by both 3 and 19

ganeshie8 (ganeshie8):

Since 3 and 19 are primes we can conclude that a^18 - 1 is also divisible by their product 3*19